presence of phenopthilein indicator 25 ml (Na2CO3+NaHCO3) mixture solution HCL bureet reading 12.8ml and his dimention 1.04(N/10).then calculate dimention of Na2CO3 per liter
1
Expert's answer
2020-09-17T07:17:51-0400

Na2CO3+HCl→NaCl+NaHCO3     

titration completion pH = 8.31  


NaHCO3+ HCl→NaCl + H2O+CO2 

titration completion pH = 3.84


Indicator transition intervals:

phenolphthalein 8.2 - 10.0

when titrating the mixture in the presence of phenolphthalein, only Na2CO3 interacts with a solution of hydrochloric acid


(25-12.8)mlN/10HCl=1/2Na2CO3

17.2mlN/10HCl=1/2Na2CO3

34.4mlN/10Na2CO3


Amount ofNa2CO3=53*34.4/34.4*1000=0.053


This amount of Na2CO3 is present in 25 mL of mixture.


This amount present in one litre of mixture

0.053/25*1000=2.12 g